OpenStudy (anonymous):

Find the sum of series (n^2 / (n+!)) x^2 ?

OpenStudy (anonymous):

\[\sum_{n=1}^{Infinity} \frac {n^2} {n+1} x^2\]

OpenStudy (anonymous):

$$\sum_{n=1}^\infty\frac{n^2}{n+1}x^2$$Well, observe:$$\frac{n^2}{n+1}=\frac{n^2-1+1}{n+1}=\frac{n^2-1}{n+1}+\frac1{n+1}=n+1+\frac1{n+1}$$Since \(n\to\infty\) this sum cannot converge...

OpenStudy (anonymous):

It converges for abs(x) < 1

OpenStudy (anonymous):

Wolfram is not always correct, for example take x = 0.

OpenStudy (anonymous):

Obviously for \(x=0\) the entire summation vanishes... but it does not converge for \(x\ne0\)... I thought you would be able to piece that together yourself...

OpenStudy (anonymous):

\(x^2\) is invariant and does not change in the summation because it does not depend on \(n\) -- thus we can pull it out:$$x^2\sum_{n=0}^\infty\left(1+n+\frac1{n+1}\right)$$

OpenStudy (anonymous):

oops, \(n-1+\frac1{n+1}\) is our summand -- messed up the sign on \(1\)

OpenStudy (anonymous):

But I've found the area of convergence using Dirichlet formula \[\lim_{n \rightarrow \infty} \left| \frac {a_{n+1}} {a_{n}} \right|=1\] which means that it should converge for \[\left| x \right|<1\] since it diverges for x = 1 and x = -1 (and of course \[\left| x \right| > 1\])

OpenStudy (anonymous):

Uh... what you've used is known as the ratio test: http://en.wikipedia.org/wiki/Ratio_test as you see, \(L=1\) is inconclusive

OpenStudy (jhannybean):

Oh. The Ratio Test.

OpenStudy (anonymous):

Consider the harmonic series \(\sum\limits_{n=1}^\infty\frac1n\):$$\lim_{n\to\infty}\left|\frac{n}{n+1}\right|=\lim_{n\to\infty}\frac{n}{n+1}=\lim_{n\to\infty}\left(1-\frac1{n+1}\right)=1$$yet we already know said series diverges!

OpenStudy (jhannybean):

the p-series also proves that the harmonic series diverges. the power of n is = 1, \(\ p \le 1\) and therefore divergent toooo.

OpenStudy (anonymous):

Oh, you're right, I overlooked that. Then the answer is that it converges only for x = 0 and that the sum is 0, right?

OpenStudy (anonymous):

Well for \(x=0\) then it converges to \(0\) yes since it's just \(0+0+0+\dots\). For any \(x\ne0\) there's no way! Be careful that you're not thinking of \(x^{2n}\) or \(x^2\) rather than \(x^2\) -- P.S. I only checked Wolfram after I answered myself.

OpenStudy (jhannybean):

What kind of series is this btw...

OpenStudy (anonymous):

I don't doubt that, it's only that I always try everything I know before I post, thanks :)

OpenStudy (anonymous):

@Jhannybean just a weird "series"... it's not a power series

OpenStudy (jhannybean):

Ohh ok. I thought this was a power series for a minute.

OpenStudy (anonymous):

I'd say it's just functional series.

OpenStudy (anonymous):

I mean technically you could consider it one if the coefficient converged to a constant value where the coefficients of all powers other than \(x^2\) are \(0\).